### v1

# Beyond Benchmark: LLMs Evaluation with an Anthropomorphic and Value-oriented Roadmap

Abstract
For Large Language Models (LLMs), a disconnect persists between benchmark performance and real-world utility. Current evaluation frameworks remain fragmented, prioritizing technical metrics while neglecting holistic assessment for deployment. This survey introduces an anthropomorphic evaluation paradigm through the lens of human intelligence, proposing a novel three-dimensional taxonomy: Intelligence Quotient (IQ)-General Intelligence for foundational capacity, Emotional Quotient (EQ)-Alignment Ability for value-based interactions, and Professional Quotient (PQ)-Professional Expertise for specialized proficiency. For practical value, we pioneer a Value-oriented Evaluation (VQ) framework assessing economic viability, social impact, ethical alignment, and environmental sustainability. Our modular architecture integrates six components with an implementation roadmap. Through analysis of 200+ benchmarks, we identify key challenges including dynamic assessment needs and interpretability gaps. It provides actionable guidance for developing LLMs that are technically proficient, contextually relevant, and ethically sound. We maintain a curated repository of open-source evaluation resources at: https://github.com/onejune2018/Awesome-LLM-Eval.
Impact Statement—As LLMs rapidly transition from research prototypes to real-world applications, the field faces a fundamental disconnect between benchmark performance and practical utility. Current evaluation practices remain fragmented, prioritizing isolated technical metrics while neglecting the developmental trajectory of LLM capabilities and their broader societal implications. This review addresses this critical gap by introducing an anthropomorphic evaluation paradigm that maps LLM assessment to human cognitive progression, through a novel four-dimensional IQ-EQ-PQ-VQ taxonomy. Crucially, our framework establishes the first comprehensive roadmap that reveals how evaluation dimensions correspond to LLMs’ developmental stages: IQ (pre-training knowledge acquisition), PQ (supervised fine-tuning expertise), EQ (reinforcement alignment), and VQ (value-oriented impact). This work provides not merely an assessment tool but a strategic compass for navigating the rapidly evolving landscape of AI evaluation. The roadmap enables stakeholders to anticipate future challenges while selecting context-appropriate evaluation strategies across the model lifecycle.
{IEEEkeywords}
Large Language Models, Evaluation, Benchmark.
1 Introduction
[…]
The quest to understand intelligence, particularly human intelligence, has been a long-standing pursuit. Throughout history, humans have employed various methods to measure and evaluate cognitive abilities, from traditional IQ tests and cognitive games to more complex assessments through education and professional achievements. This ongoing exploration aims to define, assess, and expand the boundaries of human intellect [1]. Contemporarily, the rise of machine intelligence, especially LLMs within natural language processing (NLP), has introduced a new dimension to this inquiry [2, 3]. These LLMs show remarkable capabilities in understanding and generating language, thereby prompting a critical need for effective measures and evaluation frameworks to gauge their level with respect to human intelligence [4, 5]. Formerly, the NLP community relied on simple benchmark tests to evaluate language models, focusing primarily on aspects like grammar and vocabulary. As the field progressed, more sophisticated benchmarks emerged, such as the MUC evaluations [6], which concentrated on information extraction. With the advent of deep learning, the landscape further evolved, incorporating comprehensive benchmarks like SNLI [7], SQuAD [8] and DROP [9], which not only assessed performance but also provided substantial training data.
Particularly, the emergence of large-scale pre-trained language models, such as BERT [2], marked a paradigm shift, necessitating the development of new evaluation methodologies. This led to a proliferation of shared tasks and challenges, including SemEval [10], CoNLL [11], GLUE [12], SuperGLUE [13], and XNLI [14]. These initiatives facilitated a holistic assessment of model performance, fostering continuous improvement in evaluation techniques.
As LLMs have grown in size and capability, they have demonstrated impressive performance in both zero-shot and few-shot scenarios, often rivaling fine-tuned models [1]. This has led to a transition from task-specific benchmarks to more general capability assessments, blurring the lines between distinct downstream applications. The rising benchmarks are designed to evaluate a wide range of abilities without relying on extensive training data, thus providing a more comprehensive evaluation under limited-shot conditions [15, 16, 12].
There is a need for rigorous and multifaceted evaluations not only assessing the capabilities but also ensuring alignment with human values and preferences. Pinpointing the limitations in existing evaluation techniques and devising approaches to overcome these hurdles is crucial. Nevertheless, the evaluation of LLMs is a multifaceted and resource - demanding endeavor, encompassing numerous dimensions and facets. Several recent reviews [17, 18] have examined the assessment of LLMs, yet their focus has been predominantly on benchmark tasks, datasets, and evaluation metrics, with a lack of in-depth investigation. Such an omission may compromise the validity of the evaluation process, as it overlooks crucial aspects such as practical applicability and interpretability. In an effort to bridge this gap, this paper integrates practical discourse to tackle the foundational challenges and limitations inherent in LLM evaluations that arise from varied evaluation configurations.
[…]
leaf/.style= my-box, minimum height=1.5em, fill=macaron-leaf, text=black, align=center, font=, inner xsep=2pt, inner ysep=4pt, line width=0.8pt, afterthought=, ,
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray,line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, , where level=0fill=none,draw=none,font=,rotate=90,anchor=north, where level=1fill=macaron-level1,font=, where level=2fill=macaron-level2, where level=3fill=macaron-level3, where level=4fill=macaron-level4, where level=5fill=macaron-level5, where level=1text width=16em,font=, where level=2text width=16em,font=, where level=3text width=19em,font=, where level=4text width=22em,font=, where level=5text width=18em,font=, [ [ 1. Introduction ] [ 2. LLM Evaluation Harness
/ Engineering
[ 2.1 Prominent LLM
Evaluation Harness, leaf [ Task Support, leaf ] [ Diversity of Objectives, leaf ] [ Usability, leaf ] [ Community Activity, leaf ] ] [ 2.2 Practical Guide of Modular
LLM Evaluation System, leaf [ Benchmark or Dataset Hub, leaf ] [ Model Hub, leaf ] [ Prompting Module, leaf ] [ Metrics Module, leaf ] [ Tasks Module, leaf ] [ Leaderboards and Arena Module, leaf ] [ Analysis Module, leaf ] ] ] [ 3. Anthropomorphic Evaluation
Taxonomy: IQ PQ and EQ
[ 3.1 Intelligence Quotient (IQ)
General Intelligence Evaluation, leaf ] [ 3.2 Professional Quotient (PQ)
Professional Expertise Evaluation, leaf [ Healthcare Domain, leaf ] [ Financial Domain, leaf ] [ Legal Domain, leaf ] [ Telecommunications Domain, leaf ] [ Coding Domain, leaf ] [ Software Domain, leaf ] [ Science Domain, leaf ] ] [ 3.3 Emotional Quotient (EQ)
Alignment Ability Evaluation, leaf ] ] [ 4. Value-Oriented Evaluation
(VQ - Value Quotient)
[ 4.1 Economic Value, leaf [ Cost-Benefit Ratio, leaf ] [ Return on Investment, leaf ] [ Productivity Improvement, leaf ] [ Market Acceptance, leaf ] ] [ 4.2 Social Value, leaf [ User Satisfaction, leaf ] [ Knowledge Dissemination Efficiency, leaf ] [ Public Service Improvement, leaf ] [ Education Quality Improvement, leaf ] ] [ 4.3 Ethical Value, leaf [ Fairness, leaf ] [ Transparency, leaf ] [ Privacy Protection, leaf ] [ Bias Detection, leaf ] ] [ 4.4 Environmental Value, leaf [ Energy Efficiency, leaf ] [ Carbon Footprint, leaf ] [ Sustainability, leaf ] ] ] [ 5. LLM System or
Application Evaluation
[ 5.1 RAG, leaf ] [ 5.2 Agent, leaf ] [ 5.3 Chatbot, leaf ] ] [ 6. Challenges & Future Perspectives [ 6.1 Enhanced Statistical Analysis, leaf ] [ 6.2 Composite Evaluation Systems, leaf ] [ 6.3 Interpretability and Explainability, leaf ] [ 6.4 User-Centric Benchmark, leaf ] [ 6.5 Human in the Loop Evaluation, leaf ] [ 6.6 Analytical Failure Exploration, leaf ] [ 6.7 Dynamic and Agentic Evaluation, leaf ] [ 6.8 Reproducibility Reliability Robustness, leaf ] ] [ 7. Conclusions ] ]
Figure 1: Overview of contents of this paper (zoom in).
Recent efforts have proposed taxonomies for evaluating LLMs. Specifically, [17] categorizes evaluations into knowledge, alignment, and safety, and [19] focuses on general taxonomies that prioritize abstract categorization, these frameworks often lack granularity in addressing domain-specific proficiency and human-centric practicality. Noteworthily, [20] reveal that the sampling mechanism of LLMs in decision-making exhibits a descriptive and prescriptive pattern akin to human. This enlightens an anthropomorphic perspective, allowing for a more intuitive and comprehensive assessment across scenarios.
As a potential road map to address these limitations, we observe a profound correspondence between LLM evaluation dimensions and the model’s developmental trajectory that mirrors human cognitive progression. As shown in Fig. 2, the proposed anthropomorphic framework naturally emerges from the three-stage training paradigm defining modern LLM development: Intelligence Quotient (IQ)-General Intelligence: Corresponds to capabilities developed during pre-training, where models acquire foundational knowledge through self-supervised learning on massive corpora. IQ quantifies reasoning ability and world knowledge breadth, analogous to human cognitive foundations. Professional Quotient (PQ)-Professional Expertise: Emerges from supervised fine-tuning (SFT), where models develop task-specific proficiency through instruction-response pairs. PQ measures specialized capabilities across diverse application domains. Emotional Quotient (EQ)-Alignment Ability: Cultivated through post-training reinforcement learning (RL), where models learn to align outputs with human values. EQ assesses emotional and ethical resonance with human preferences beyond mere task completion. Unlike broad ’knowledge’ dimension in [19], our IQ evaluation explicitly quantifies foundational reasoning and world knowledge breadth, while PQ introduces a structured evaluation of task-specific expertise, which existing frameworks neglect. Furthermore, EQ extends beyond [17]’s safety-centric alignment to encompass emotional and ethical alignment with human values, ensuring outputs resonate with user preferences and societal norms.
Early stages of LLM evaluation mainly focus on IQ, ensuring that the models had a broad base of world knowledge. As pre-training techniques and data engineering matured, the emphasis shifted to PQ, evaluating the model’s ability to solve specific practical tasks. Now, as models become proficient in these tasks, EQ has become increasingly important. For IQ and PQ, there are well-established benchmarks such as MMLU [16], GPQA [21], MATHQA[22] for IQ, and HumanEval [23], IFEval [24] for domain-specific PQ. For EQ, while there are no strict benchmarks, tools like Alignbench [25], MT-Bench [26], and Arena-Hard [27] provide some coverage, though they often use third-party AI as evaluators, making them more aligned with AI preferences than human preferences.
Figure 2: The proposed technical evolutionary tree of the LLM evaluation, following the structure in [28] for RAG. The anthropomorphic evaluation framework: IQ-EQ-PQ taxonomy with evolutionary correspondence to LLM training stages. Intelligence Quotient (IQ)-General Intelligence denotes knowledge capacity acquired by pre-training, reflecting foundational reasoning and world knowledge breadth. Professional Quotient (PQ)-Professional Expertise represents task capability developed through supervised fine-tuning (SFT), measuring proficiency in specialized domains. Emotional Quotient (EQ)-Alignment Ability represents human preference alignment achieved through RL post-training, encompassing emotional and ethical resonance with human values.
As depicted in Fig. 1, this review transcends conventional LLM evaluation paradigms by introducing a transformative framework, that bridges the critical gap between technical performance metrics and real-world societal impact. We pioneer an anthropomorphic evaluation taxonomy that fundamentally reimagines how we assess LLM capabilities, moving beyond fragmented benchmarks toward a holistic roadmap understanding of AI intelligence. Our work establishes the first comprehensive bridge between machine cognition and human-centric value systems, positioning the evaluation not merely as a technical exercise but as a crucial determinant of responsible AI deployment. The contributions of this paper are as follows:
  • Revolutionizing LLM Evaluation Taxonomy: We present the first systematic engineering framework that transcends traditional categorization approaches, offering a granular analysis of over 200 evaluation benchmarks/frameworks across six dimensions. Our taxonomic structure not only maps the current landscape with unprecedented precision, but also reveals hidden interconnections between seemingly disparate evaluation techniques, exposing critical gaps that have hindered the development of truly comprehensive assessment protocols.
  • Anthropomorphic Intelligence Framework: Breaking free from the limitations of single-dimensional evaluations, we introduce a paradigm-shifting anthropomorphic framework that conceptualizes LLM capabilities through the lens of human intelligence. Our tripartite IQ-EQ-PQ model (Intelligence Quotient, Emotional Quotient, and Professional Quotient) represents the first holistic approach (to our best knowledge) that simultaneously captures what LLMs know, how they apply knowledge, and why their outputs resonate with human values. This framework transforms evaluation from a technical checklist into a meaningful assessment of AI’s alignment with human cognitive and social structures.
  • Pioneering Value-Oriented Evaluation (VQ): We establish the foundational principles for Value Quotient (VQ) assessment—the first systematic methodology to quantify LLMs’ broader societal impact beyond technical metrics. By integrating economic viability, ethical alignment, social responsibility, and environmental sustainability into a unified evaluation framework, we shift the discourse from "can it work?" to "should it work?" and "how does it benefit society?" This represents a critical evolution from capability-focused assessment to value-driven evaluation.
  • Practical Implementation Blueprint: Beyond theoretical constructs, we deliver an actionable, step-by-step and modular evaluation system that bridges the chasm between academic research and industrial deployment. Our evaluation framework addresses the critical disconnect between benchmark performance and real-world functionality, providing concrete strategies for evaluating LLMs within complex application ecosystems (RAG systems, agents, chatbots) while accounting for the full lifecycle of model deployment and maintenance.
  • Future-Proof Evaluation Roadmap: We articulate a six-tiered evolutionary path for LLM evaluation that anticipates the field’s trajectory over the next decade. This forward-looking perspective identifies not just current limitations but also the emerging challenges at the intersection of statistical rigor, interpretability, user experience, system reliability, dynamic adaptation, and value creation—providing a strategic compass for navigating the rapidly evolving landscape of AI assessment.
2 LLM Evaluation Harness / Engineering
Table 1: Comparison of LLM Evaluation Harnesses or Toolkits, IF denotes Instruction Following.
Toolkit Ease of Use Modularity Explainability Metrics Richness Multi-Task Efficiency Testing IF
Openbench(2025.8) *** *** ** ** ** ** No
Eval-assist(2025.2) *** ** ** ** ** ** No
Evalchemy(2025.1) *** ** ** *** *** *** Yes
Evalscope(2024.12) *** *** *** *** ** *** Yes
LeaderboardFinder(2024.9) ** ** ** ** ** * No
Vertex AI Studio(2024.7) ** *** ** ** ** ** No
LLMeBench(2024.6) *** *** ** *** ** * No
LightEval(2024.5) *** *** *** *** ** * No
Athina Evals(2024.4) *** ** ** ** ** * No
Prometheus Eval(2024.3) *** ** ** ** ** * No
LLM Comparator(2024.2) *** ** *** *** ** * No
Azure AI Studio(2024.2) *** ** ** ** ** ** No
Uptrain(2024.2) *** ** *** *** ** *** No
Evidently(2024.1) *** *** ** *** ** ** No
LM Evaluation Harness(2023.12) ** ** ** ** ** * No
EVAL(2023.11) ** ** ** ** ** * No
AutoEvals(2023.10) *** ** ** ** ** * No
LLM Benchmarker Suite(2023.9) ** ** ** ** ** * No
Arthur Bench(2023.8) *** ** *** *** ** * Yes
OpenCompass(2023.8) *** *** ** *** *** * No
DeepEval(2023.8) ** *** *** *** ** * Yes
CONNER(2023.8) ** ** ** ** ** * No
Amazon Bedrock(2023.7) *** ** ** *** *** *** No
Alpaca Eval(2023.7) ** ** ** ** ** * No
h2o-LLM-eval(2023.7) *** ** *** ** ** * No
Parea AI(2023.6) *** *** ** *** *** * No
Prompt Flow(2023.6) *** ** * ** ** * Yes
TruLens(2023.6) ** ** *** *** ** * Yes
LangSmith(2023.5) ** ** *** *** ** * Yes
SuperCLUE(2023.5) ** * * ** * * No
PandaLM(2023.4) *** ** ** ** ** * No
HELM(2023.3) ** ** ** ** ** * No
Auto-Evaluator(2023.2) *** ** ** ** ** * Yes
LM Evaluation(2023.1) ** ** ** ** ** * No
FlagEval(2022.12) ** *** ** ** ** * No
Weights & Biases(2022.7) *** *** ** *** *** * No
2.1 Prominent LLM Evaluation Harness
Table 1 summarizes a range of prominent LLM evaluation tools and frameworks, each representing different organizations’ and individuals’ efforts to enhance assessment methodologies [29, 30]. By analyzing these tools, we can better understand their strengths and limitations, providing recommendations for future improvements and deployments. The comprehensive analysis of these LLM evaluation harnesses reveals a spectrum of strengths and weaknesses across several critical dimensions. When it comes to ease of use, some platforms like OpenCompass and Azure AI Studio stand out for their user-friendly interfaces, streamlining the process for both novice and experienced researchers. However, the modularity of these tools varies; FlagEval and Weights & Biases offer high levels of customization, allowing users to integrate specific components as needed, which is particularly beneficial for complex or specialized research projects.
Explainability, with Arthur Bench and LangSmith provides robust mechanisms to interpret model behavior, an essential aspect for ensuring transparency and trust in AI systems. In terms of reproducibility, most of the listed tools, including SuperCLUE and DeepEval, ensure that experiments can be reliably replicated, which is fundamental for the scientific method. The open-source nature of many of these tools, such as DeepEval and Parea AI, fosters a collaborative environment.
The richness of the metrics provided by these evaluation harnesses is also noteworthy. While some, like Arthur Bench and TruLens, offer a wide array of detailed performance indicators, others may focus on a more limited but still informative set. Multi-task support is another area where there’s a significant difference, with Azure AI Studio and Amazon Bedrock excelling in handling a broad range of tasks, from natural language understanding to generation, thereby providing a more holistic assessment of LLMs.
Speed and efficiency testing are crucial for practical applications, yet not all toolkits include this feature. Tools like Azure AI Studio and Vertex AI Studio incorporate speed and resource efficiency evaluations, which are vital for real-world deployment considerations. Lastly, the ability to assess alignment and instruction following, an increasingly important aspect of LLMs, is present in select platforms, such as Arthur Bench and Prompt Flow, which provide insights into how well models adhere to human values and follow specific instructions, a critical consideration for safe and effective AI.
Overall, the landscape of LLM evaluation harnesses is diverse, with each tool offering a different balance of features. Researchers and developers must carefully consider their specific needs and the characteristics of the available tools when selecting the most appropriate one for their work. By leveraging the strengths of these platforms, the field can continue to advance the quality, reliability, and applicability of LLMs, contributing to the broader goals of artificial intelligence.
Task Support and Diversity of Objectives
The existing evaluation tools cover a wide array of tasks, including direct assessment, pairwise ranking, question-answering, summarization, translation, and code generation. This diversity reflects the complexity and variability of real-world applications. For instance, Arthur Bench supports multiple task types, such as QA, summarization, and translation, making it a versatile evaluation platform. Additionally, many tools allow users to customize tasks, for instance, athina-evals and PandaLM, which is valuable for specific research or industrial applications.
Usability and Community Activity
Usability is a critical factor in determining the widespread adoption of an evaluation tool. The table indicates that most tools have achieved high standards of usability, with intuitive interfaces and documentation. For example, LightEval and autoevals are noted for their high usability, providing straightforward access for users. Some tools also integrate automated processes to further simplify the evaluation workflow. Prompt flow by Microsoft, for instance, aims to enhance product quality through simplified development processes.
Community activity is another key indicator. High activity typically means continuous support and updates, along with a strong user base contributing feedback and improvements. Projects like EVAL (OpenAI) and lm-evaluation-harness (EleutherAI) exhibit strong community engagement, which not only drives the iterative improvement of the tools but also provides a wealth of resources and support for users.
Figure 3: Typology of the LLM Evaluation Modules.
2.2 Implementation Roadmap of Modular Evaluation System
A modular LLM evaluation framework or harness in general consists of: benchmark or dataset hub, model hub, prompting module, metrics, monitoring and experiment management, arena or leaderboard (as shown in Fig. 3).
The evaluation framework leverages distinct modules, delineating three primary paradigms: metrics-centered assessment, human-centered assessment (Human Judgment), and model-centered peer review (LLMs as Evaluators). In the metrics-centered assessment paradigm, task-specific performance indicators—such as F1 score, Exact Match, and Perplexity [31]—are commonly employed to ascertain the accuracy of generated outputs, particularly in classification-oriented tasks. The human-centered assessment approach emphasizes human’s qualitative analysis of LLM-generated content, focusing on attributes like clarity, coherence, and factual correctness [32]. Notably, there has been a surge in interest towards human evaluations utilizing the Elo rating system [33], which offers a structured methodology for comparative assessment. Since human evaluations are time-consuming, using model-centered peer review (LLMs as Evaluators) has become a popular alternative for assessing model performance. [34].
2.2.1 Benchmark or Dataset Hub
It is crucial to select appropriate benchmark datasets that accurately reflect the models’ capabilities. Analogous to human intelligence, LLM abilities can be classified into three interrelated dimensions: General Intelligence (IQ, Intelligence Quotient), Alignment Ability (EQ, Emotional Quotient), and Professional Expertise (PQ, Professional Quotient). Correspondingly, benchmark datasets are categorized into general capability benchmarks, alignment benchmarks, and domain-specific benchmarks. General capability benchmarks serve as foundational assessments, often employed at the time of an LLM’s release to gauge its broad-spectrum performance (e.g., MMLU [35], HumanEval [36]). Domain-specific benchmarks focus on specialized areas, evaluating LLMs’ proficiency in particular fields such as telecommunications with TeleQnA [37]. Furthermore, alignment benchmarks scrutinize LLMs’ adherence to diverse tasks and ethical guidelines, exemplified by AlignBench [25]. Additional benchmarks like FOFO [38] assess specific competencies, such as format-following capabilities. Detailed descriptions of each category are provided in Section 3.
2.2.2 Model Hub
This section provides insights into various models, ensuring a fair evaluation by mitigating risks such as data contamination and avoiding biased comparisons. It addresses considerations for selecting models based on their training methodologies, access to external resources, and fine-tuning on specific benchmarks versus pre-training only.
2.2.3 Prompting Module
After selecting suitable benchmarks and models, the subsequent step involves designing prompts and configuring decoding parameters for response generation. In the prompt design phase, decisions are made regarding the type of prompting strategy—whether zero-shot, few-shot, or chain-of-thought—to employ. The configuration of decoding parameters, including temperature settings, plays a critical role in optimizing model output. Proper setup ensures that the evaluation not only tests the LLM’s inherent capabilities but also its adaptability under varying conditions.
2.2.4 Metrics Module
The evaluation of LLMs necessitates the selection of appropriate metrics that align with specific applications and intended use cases. Given the broad spectrum of LLM applications, from machine translation and text summarization to conversational agents, the choice of evaluation metrics would be beneficial to not only reflect technical performance, but also be closely tied to business needs and application contexts. An effective evaluation framework allows researchers and developers to gain deep insights into the strengths and limitations of LLMs, guiding further improvements and optimizations. The evaluation requires a dual focus on technical performance and business impact. Technical metrics assess the model’s linguistic and functional capabilities, while business metrics measure user engagement, operational efficiency, and cost-effectiveness.
(1) Technical Metrics: The choice of metrics would be closely aligned with the application. For instance, in machine translation, where the goal is to generate translations that are both accurate and fluent, metrics such as BLEU [39] and METEOR [40] have been widely adopted. These token overlap-based metrics measure the n-gram overlap between the generated text and the reference, providing an indication of how well the model’s output matches human-generated translations. In contrast, for tasks like sentiment analysis, precision, recall, and F1 score become more relevant, as they focus on the model’s ability to correctly classify the sentiment of a given text.
Considering the diverse range of LLM applications—from machine translation and summarization to dialogue systems and code generation—it is essential to adopt a multi-layered evaluation framework that reflects the linguistic phenomena at play. Table 2 presents a taxonomy of technical metrics for LLM evaluation, organized into five broad levels: (1) Lexical and Morphological, (2) Syntactic, (3) Semantic, (4) Pragmatic and Discourse, and (5) Factuality and Explainability.
Lexical and Morphological Metrics: These metrics focus on token- or character-level correspondence and morphological variation. Traditional n-gram overlap measures such as BLEU [39] and ROUGE-N/L [41] quantify the proportion of exact contiguous matches between hypothesis and reference, while edit-distance scores like Translation Error Rate (TER) gauge the minimal sequence of insertions, deletions, and substitutions required to transform one string into another. Complementing these, word-order distances—including RIBES [42] and Kendall’s $\tau$ —penalize token reordering, providing insight into the impact of syntactic shifts on surface similarity. For tasks sensitive to finer-grained discrepancies, character error rate (CER) and word error rate (WER) compute the frequencies of low-level insertion, deletion, and substitution errors, as commonly used in ASR and OCR evaluation. Subword-overlap metrics such as chrF [43] and BPE-F1 further refine this analysis by measuring similarity over character n-grams or byte-pair encoded segments, thus capturing partial matches that evade pure token-level statistics.
Syntactic Metrics: it assesses the preservation of grammatical structure and targeted syntactic phenomena. Constituency and dependency parse-tree matching metrics—PARSEVAL [44] precision, recall, and F1 for bracket structures, alongside Unlabeled and Labeled Attachment Scores (UAS/LAS) [45] for dependency relations—offer a principled basis for comparing predicted and gold parses. To probe a model’s command of specific constructions, targeted syntactic evaluation frameworks such as Targeted Syntactic Evaluation (TSE) [46] deploy minimal-pair sentences to test capabilities like subject–verb agreement, while syntactic tree-edit distance measures the minimal sequence of tree operations to align two parse trees, yielding a granular account of structural divergence.
Semantic Metrics: At semantic level, evaluation emphasizes meaning preservation, inference, and fidelity. Embedding-based approaches—BERTScore [47], MoverScore [48], BLEURT [49], COMET [50], and similar methods—leverage contextualized vectors extracted from pre-trained models to compute cosine similarities or Earth Mover’s distances, thus capturing nuanced semantic alignment between hypothesis and reference. Entailment-driven metrics such as the Document-Aware Entailment model (DAE) [51] treats the generation task as a natural language inference problem, classifying whether outputs are entailed by, neutral to, or contradictory with source texts. Question-answering frameworks, including QuestEval [52] and QAFactEval [53], automatically generate questions from the source or candidate summary and compare model-predicted answers to gold responses, thereby quantifying semantic fidelity via answer accuracy. Specialized LLM-based judges—either prompt-based few-shot evaluators or fine-tuned discrimination models—have emerged as learnable arbiters of output quality, scoring on dimensions such as factuality, coherence, and naturalness [54].
Pragmatic and Discourse Metrics: The metrics in this part capture coherence, cohesion, style, and diversity across larger textual spans. Entity Grid models [55] track the distribution and syntactic roles of discourse entities across sentences to quantify thematic coherence, while Rhetorical Structure Theory (RST) tree comparisons [56] evaluate whether logical and rhetorical relations are preserved. Readability and stylistic consistency are measured by indices such as Flesch–Kincaid readability tests [57] combine sentence length and word complexity into difficulty scores, alongside formality and sentiment metrics [58] that assess register and affective tone. To detect degeneracy and encourage lexical variety, diversity measures like distinct-n [59] compute the proportion of unique n-grams in the generated text, whereas [60] quantify the recurrence of identical n-grams within and across sentences.
Factuality & Explainability Metrics: Factual consistency metrics such as FactCC [61] verify whether key propositions in the generated output align with source material. Calibration and uncertainty metrics, including Expected Calibration Error (ECE) and Maximum Calibration Error (MCE), measure discrepancies between predicted probabilities and observed accuracy, while entropy-based measures of predictive and semantic uncertainty signal where the model is least confident.
By selecting and combining technical metrics, we can gain a deeper understanding of models’ strengths and weaknesses, leading to more informed decisions in the development and deployment of LLMs. Future work may focus on developing more sophisticated and context-aware metrics that can better capture the nuances of natural language, thus bridging the gap between automatic evaluations and human judgment.
Table 2: Basic Metrics Taxonomy for LLM Evaluation (Technical vs. Business).
Dimension Category Metric Description / Use-case
Technical Lexical & Morphological BLEU [39] Precision-based n-gram overlap (0–1); $\uparrow$ is better.
ROUGE-N/L [41] Recall-oriented n-gram (N) and LCS-based (L) overlap.
METEOR [40] Unigram alignment with synonymy/stemming; harmonic mean.
TER Translation Edit Rate via edit distance.
chrF [43] Character n-gram F-score for finer-grained matching.
Syntactic PARSEVAL [44] Constituency precision/recall/F1 on parse trees.
UAS / LAS [45] Unlabeled / Labeled dependency attachment scores.
TSE [46] Targeted syntactic evaluation via minimal-pair sentences.
Semantic BERTScore [47] Contextual-embedding cosine similarity (BERT/RoBERTa).
MoverScore [48] Earth Mover’s Distance on contextual embeddings.
COMET [50] Learned metric using cross-lingual embeddings.
QuestEval [52] QA-based semantic fidelity assessment.
Pragmatics & Discourse Entity Grid [55] Entity transition coherence modeling.
distinct-n [59] Lexical diversity via unique n-gram ratio.
Flesch-Kincaid [57] Readability via sentence/word complexity.
Factuality & Explainability FactCC [61] Factual consistency via source alignment.
ECE / MCE [62] Expected / Maximum calibration error for confidence.
BLANC [63] Reference-less metric using masked LM.
Business User Engagement Visited Count of unique users accessing the LLM interface.
Submitted Ratio of users who submit prompts vs. total visitors.
Responded Proportion of error-free system outputs delivered.
Viewed Frequency of users viewing generated responses.
Clicks Number of reference-document clicks from outputs.
Interaction User acceptance rate Context-specific adoption (e.g., thumbs-up, text reuse).
LLM conversation Mean dialogue sessions per user.
Active days Distinct days each user interacts with the LLM.
Interaction timing Avg. prompt-to-response latency + dwell time.
Response Quality Prompt / response length Avg. tokens in queries and replies.
Edit distance Textual delta between prompt and generated output.
Feedback & Retention User feedback Volume of up/down votes or explicit ratings.
DAU / WAU / MAU Daily/Weekly/Monthly Active Users.
User return rate % of prior-period users who return.
Performance Requests per second Peak sustained throughput (concurrency).
Tokens per second Streaming generation speed.
Time to first token Latency p50/p95 from query to first byte.
Error rate Fraction of failed requests (auth, rate-limit, etc.).
Reliability Success-to-total request ratio.
Latency End-to-end response time (avg / p95 / p99).
Cost GPU / CPU utilization Resource efficiency (tokens per GPU-hour).
LLM API cost Third-party token or query charges.
Infrastructure cost Storage, bandwidth, compute amortization.
Operation cost Maintenance, security, support staff spend.
(2) Business Metrics: Evaluating a system’s performance and impact on business is multifaceted. The metrics used to gauge the success of an LLM application can be broadly categorized into several key areas: user engagement and utility, user interaction, quality of response, user feedback and retention, performance, and cost. Each category provides insights into the operational efficiency and user experience.
User Engagement and Utility Metrics: Both are fundamental in assessing the initial attractiveness and usability of an LLM application. These metrics include the number of users who visited the LLM app feature, submitted prompts, received responses without errors, viewed responses, and clicked on reference documentation provided by the LLM. A high rate of visits and submissions indicates a strong interest and active use of the LLM, while the absence of errors and the viewing of responses suggest that the LLM is providing value to its users. Clicks on reference documentation can also indicate that the LLM is effectively guiding users towards additional resources, enhancing their overall experience.
User Interaction Metrics: These metrics delve deeper into how users engage with the LLM over time. The frequency of user acceptance, the average number of LLM conversations per user, the number of active days using LLM features, and the average interaction timing all provide a comprehensive view of user behavior. For instance, a higher user acceptance rate, especially in conversational scenarios, suggests that the LLM is meeting or exceeding user expectations. Monitoring the average number of conversations and active days can help identify power users and potential areas for improvement. Interaction timing, including the latency between prompts and responses, is crucial for ensuring that the LLM remains responsive and engaging.
Response Quality Metrics : This is paramount for maintaining user trust and satisfaction. Average lengths of prompts and responses, as well as edit distance metrics, offer quantitative measures of the LLM’s ability to generate coherent and relevant content. Edit distance metrics, in particular, can serve as an indicator of the degree of customization and refinement in the LLM’s output, reflecting its adaptability to user needs. High-quality responses not only improve user experience but also contribute to the LLM’s reputation and credibility.
Feedback and Retention Metrics: Direct feedback like thumbs up/down ratings, is invaluable for understanding user sentiment and making data-driven improvements. Additionally, tracking daily, weekly, and monthly active users, along with user return rate, helps in assessing the stickiness of the LLM application. A high return rate indicates that the LLM is delivering consistent value, encouraging users to continue using. Analyzing these metrics can guide the development of strategies to enhance user retention and satisfaction.
Performance Metrics: Performance metrics are essential for ensuring that the LLM operates efficiently and reliably. As supported in LLMPerf, key performance indicators include requests per second (concurrency), tokens per second, time to first token render, error rates, reliability, and latency. These metrics provide a practical overview of the LLM’s capabilities, helping to identify bottlenecks and areas for optimization. For example, a low error rate and high reliability are indicative of a robust and stable system, while minimizing latency ensures a smooth and responsive user experience.
Cost Metrics: GPU/CPU utilization, LLM calls cost, infrastructure cost, and operation cost all contribute to the total cost of ownership. By monitoring these costs, organizations can make informed decisions about resource allocation and scaling. For instance, optimizing GPU/CPU utilization can lead to cost savings, while carefully managing infrastructure and operation costs ensures LLMs economically viable.
A comprehensive evaluation of an LLM system requires a balanced approach that considers both qualitative and quantitative aspects. By leveraging the metrics outlined in Table 2, businesses can gain a holistic understanding of their LLM’s performance, enabling them to continuously refine and improve the service to meet the evolving needs of their users.
2.2.5 Tasks Module
The Tasks Module is a critical component within the evaluation framework for LLMs, designed to systematically assess model performance across a wide array of tasks. This module aims to provide a comprehensive and diverse set of challenges that can effectively evaluate various aspects of LLM capabilities, including language understanding, reasoning, and generation, etc. The selection of tasks is crucial as it directly influences the breadth and depth of the evaluation, guaranteeing that models undergo evaluation in situations closely resembling practical applications.
To achieve this goal, the Tasks Module incorporates both conventional and innovative tasks. Conventional tasks include those found in established benchmarks such as GLUE [12], which focus on natural language understanding. However, recognizing the limitations of these benchmarks, newer frameworks like BIG-bench [64] have expanded the scope to include more complex and varied challenges. These tasks are designed to push the boundaries of what LLMs can do, thereby identifying areas where further improvements are needed.
Moreover, the Tasks Module emphasizes the importance of real-world applicability. For instance, HELM [54] introduces a hierarchical categorization framework which spans 16 distinct scenarios, each represented by <task, domain, language> triples. This approach ensures that evaluations cover a broad spectrum of user-oriented tasks, from simple instructions to intricate reasoning problems. Additionally, OpenCompass [30] extends its scope beyond traditional areas like language and reasoning to encompass comprehension and subject-specific evaluations, offering a more holistic view of LLM capabilities.
The inclusion of dynamic and adaptable tasks is another hallmark of modern evaluation frameworks. FlagEval [65], for example, allows users to dynamically combine capabilities, tasks, and metrics into ternary groups, significantly enhancing the flexibility and adaptability of the evaluation process. This modular design enables researchers to tailor evaluations to specific needs or emerging trends in LLM development.
Thus, the Tasks Module serves as a cornerstone for evaluating LLMs, providing a structured yet flexible environment that can accommodate both established and new challenges. By continuously updating and refining the task set, it plays a pivotal role in advancing the SOTA in LLM technology.
2.2.6 Leaderboards and Arena Module
The Leaderboards and Arena Module represents an essential tool for benchmarking and comparing LLMs in a transparent and competitive manner. Leaderboards offer a standardized platform where models can be evaluated against predefined datasets and metrics, while Arenas introduce a more interactive approach, leveraging human preferences to rank models based on direct comparisons [66]. Together, these modules facilitate a deeper understanding of LLM performance and promote continuous improvement within the research community.
Leaderboards, such as those provided by Hugging Face’s Open LLM Leaderboard, serve as centralized repositories for sharing and comparing evaluation results. They typically highlight key datasets like ARC [67], HellaSwag [68], MMLU [16], and TruthfulQA [69], selected for their ability to challenge LLMs in different ways. By making evaluation results public, leaderboards foster transparency and encourage collaborative efforts towards improving LLM technologies.
Arenas, on the other hand, adopt a more interactive evaluation paradigm. Platforms like Chatbot Arena [66] allow users to compare outputs from multiple LLMs for a given query, using human preferences as the primary metric. The Elo scoring mechanism is employed to dynamically adjust scores based on user feedback, providing a scalable and adaptive ranking system. It not only streamlines the evaluation process but also captures nuanced differences in performance that might not be evident by automated metrics alone.
By engaging the broader community, the Arena Module enhances the relevance and reliability of evaluations, ensuring that models are judged based on their actual utility rather than just theoretical benchmarks. Furthermore, the Arena Module addresses some of the limitations inherent in static leaderboards. While leaderboards provide a snapshot of performance at a given time, arenas offer ongoing assessments that evolve with user interactions. This dynamic nature helps maintain the integrity and relevance of evaluations, reducing the risk of data leakage and ensuring that benchmarks remain challenging and informative. Therefore, the Leaderboards and Arena Module complements the Tasks Module by providing both standardized and interactive platforms for evaluating LLMs.
Table 3: 64 typical Intelligence Quotient (IQ)-General Intelligence evaluation benchmarks for LLMs.
Name Year Task Type Institution Evaluation Focus Datasets Url
MMLU-Pro [16] 2024 Multi-Choice Knowledge TIGER-AI-Lab Subtle Reasoning, Fewer Noise MMLU-Pro link
DyVal [70] 2024 Dynamic Evaluation Microsoft Data Pollution, Complexity Control DyVal link
PertEval [71] 2024 General USTC Knowledge capacity PertEval link
LV-Eval [72] 2024 Long Text QA Infinigence-AI Length Variability, Factuality 11 Subsets link
LLM-Uncertainty-Bench [73] 2024 NLP Tasks Tencent Uncertainty Quantification 5 NLP Tasks link
CommonGen-Eval [74] 2024 Generation AI2 Common Sense CommonGen-lite link
MathBench [75] 2024 Math Shanghai AI Lab Theoretical and practical problem-solving Various link
AIME [76] 2024 Math MAA American Invitational Mathematics Examination Various link
FrontierMath [77] 2024 Math Epoch AI Original, challenging mathematics problems Various link
FELM [78] 2023 Factuality HKUST Factuality 847 Questions link
Just-Eval-Instruct [79] 2023 General AI2 Mosaic Helpfulness, Explainability Various link
MLAgentBench [80] 2023 ML Research snap-stanford End-to-End ML Tasks 15 Tasks link
UltraEval [81] 2023 General OpenBMB Lightweight, Flexible, Fast Various link
FMTI [82] 2023 Transparency Stanford Model Transparency 100 Metrics link
BAMBOO [83] 2023 Long Text RUCAIBox Long Text Modeling 10 Datasets link
TRACE [84] 2023 Continuous Learning Fudan University Continuous Learning 8 Datasets link
ColossalEval [85] 2023 General Colossal-AI Unified Evaluation Various link
LLMEval² [86] 2023 General AlibabaResearch Wide and Deep Evaluation 2,553 Samples link
BigBench [87] 2023 General Google knowledge, language, reasoning Various link
LucyEval [88] 2023 General Oracle Maturity Assessment Various link
Zhujiu [89] 2023 General IACAS Comprehensive Evaluation 51 Tasks link
ChatEval [90] 2023 Chat THU-NLP Human-like Evaluation Various link
FlagEval [91] 2023 General THU Subjective and Objective Scoring Various link
Chain-of-thought [92] 2023 Reasoning UE Complex Problem Solving GSM8k, MATH link
AlpacaEval [93] 2023 General tatsu-lab Automatic Evaluation Various link
GPQA [21] 2023 General NYU Graduate-Level Google-Proof QA Various link
MuSR [94] 2023 Reasoning Zayne Sprague Narrative-Based Reasoning 756 link
FreshQA [95] 2023 knowledge FreshLLMs Current World Knowledge 599 link
AGIEval [96] 2023 general Microsoft Human-Centric Reasoning NA link
SummEdits [97] 2023 general Salesforce Inconsistency Detection 6,348 link
ScienceQA [98] 2022 Reasoning UCLA Science Reasoning 21,208 link
e-CARE [99] 2022 Reasoning HIT Explainable Causality 21,000 link
BigBench Hard [64] 2022 Reasoning BigBench Challenging Subtasks 6,500 link
PlanBench [100] 2022 Reasoning ASU Action Planning 11,113 link
MGSM [101] 2022 Math Google Grade-school math problems in 10 languages Various link
MATH [102] 2021 Math UC Berkeley Mathematical Problem Solving Various link
GSM8K [103] 2021 Math OpenAI Diverse grade school math word problems Various link
SVAMP [104] 2021 math Microsoft Arithmetic Reasoning 1,000 link
SpartQA [105] 2021 Reasoning MSU Textual Spatial QA 510 link
MLSUM [106] 2020 general Thomas Scialom News Summarization 535,062 link
Natural Questions [107] 2019 Language, Reasoning Google Search-Based QA 300,000 link
ANLI [108] 2019 Language, Reasoning Facebook AI Adversarial Reasoning 169,265 link
BoolQ [109] 2019 Language, Reasoning Google Binary QA 16,000 link
SuperGLUE [13] 2019 Language, Reasoning NYU Advanced GLUE Tasks NA link
DROP [9] 2019 Language, Reasoning UCI NLP Paragraph-Level Reasoning 96,000 link
HellaSwag [68] 2019 Language, Reasoning AI2 Commonsense Inference 59,950 link
Winogrande [110] 2019 Language, Reasoning AI2 Pronoun Disambiguation 44,000 link
PIQA [111] 2019 Language, Reasoning AI2 Physical Interaction QA 18,000 link
HotpotQA [112] 2018 Language, Reasoning HotpotQA Explainable QA 113,000 link
GLUE [12] 2018 Language, Reasoning NYU Foundational NLU Tasks NA link
OpenBookQA [113] 2018 Language, Reasoning AI2 Open Book Exams 12,000 link
SQuAD2.0 [114] 2018 Language, Reasoning Stanford University Unanswerable Questions 150,000 link
ARC [67] 2018 Language, Reasoning AI2 AI2 Reasoning Challenge 7,787 link
SWAG [115] 2018 Language, Reasoning AI2 Adversarial Commonsense 113,000 link
CommonsenseQA [116] 2018 Language, Reasoning AI2 Commonsense Reasoning 12,102 link
RACE [117] 2017 Language, Reasoning CMU Exam-Style QA 100,000 link
SciQ [118] 2017 Language, Reasoning AI2 Crowd-Sourced Science 13,700 link
TriviaQA [119] 2017 Language, Reasoning AI2 Distant Supervision 650,000 link
MultiNLI [120] 2017 Language, Reasoning NYU Cross-Genre Entailment 433,000 link
SQuAD [8] 2016 Language, Reasoning Stanford University Wikipedia-Based QA 100,000 link
LAMBADA [121] 2016 Language, Reasoning CIMEC Discourse Context 12,684 link
MS MARCO [122] 2016 Language, Reasoning Microsoft Search-Based QA 1,112,939 link
2.2.7 Analysis Module
It is designed to interpret and synthesize the extensive data generated during evaluations. This module integrates advanced analytical techniques to provide meaningful insights into model performance, thereby guiding future improvements and informing strategic decisions regarding LLM deployment. Specifically, it addresses critical areas such as Monitoring, Logs, Experiment Management, Visualization, and Statistics, each of which plays an essential role in ensuring comprehensive and actionable evaluations.
Monitoring is essential to tracking the performance of LLMs during evaluation. Continuous monitoring allows evaluators to detect anomalies or deviations from expected behavior promptly. The module employs real-time feedback mechanisms to ensure that models are performing consistently across various tasks. Monitoring also facilitates early detection of issues related to computational resources, enabling timely adjustments to optimize efficiency. Moreover, continuous monitoring supports iterative development cycles.
Logs serve as a record of interactions between LLMs and the evaluation environment, capturing inputs, outputs, and intermediate states. They are indispensable for post-hoc analysis and debugging. It also plays a role in auditing and compliance, ensuring that evaluations adhere to ethical standards and regulatory requirements. By maintaining thorough logs, the Analysis Module enhances transparency and accountability.
Experiment Management is vital for systematic evaluations. It involves defining protocols, managing datasets, and controlling variables to ensure reproducibility and comparability of results. Platforms like OpenCompass [30] offer versatile experimental settings, including zero-shot, few-shot, and Chain-of-Thought (CoT) configurations, allowing researchers to explore different facets of LLM capabilities. Effective experiment management also includes version control and documentation practices, ensuring that each experiment can be replicated or extended by other researchers.
Visualization tools transform complex evaluation data into intuitive and accessible formats, enhancing the interpretability of results. The LLM Comparator [123] provides an interactive table and visualization summary that enable users to inspect individual prompts and their responses in detail. These visual aids facilitate the identification of trends, outliers, and correlations, supporting deeper analyses. Visualization also plays a key role in communicating findings to stakeholders who may not have technical expertise, ensuring that insights from evaluations are widely understood and acted upon.
Statistical analysis. Techniques such as hypothesis testing, regression analysis, and confidence interval estimation are employed to quantify uncertainties and validate findings. Statistical rigor also helps in identifying significant factors influencing model performance, informing strategies for optimization and enhancement. By applying robust statistical practices, the Analysis Module ensures that evaluations yield accurate and trustworthy insights.
3 Anthropomorphic Evaluation: IQ, PQ, EQ
It necessitates to draw an analogy with human intelligence, categorizing their abilities into three interconnected dimensions: General Intelligence (IQ, Intelligence Quotient), Alignment Ability (EQ, Emotional Quotient), and Professional Expertise (PQ, Professional Quotient). It allows us to gain a more nuanced and easier understanding of their performance in practical scenarios. It also provides guidance to the enhancement of their cognitive, social, and professional competencies.
3.1 General Intelligence Evaluation (IQ)
General Intelligence of an LLM refers to its foundational cognitive capabilities (IQ). It encompasses the model’s ability to understand, reason, and learn from a wide array of textual data. This includes the capacity for language comprehension, logical reasoning, and the generation of coherent and contextually appropriate responses. The IQ of an LLM is analogous to the human mind’s ability to process information from various domains and to apply general knowledge flexibly. Crucially, IQ corresponds to capabilities developed during pre-training, where models acquire foundational knowledge through self-supervised learning on massive corpora, reflecting the breadth of world knowledge and reasoning ability that forms the bedrock of LLM performance.
Different benchmarks offer diverse perspectives through their unique approaches and task types (Table 3). The MMLU benchmark [35] encompasses a diverse array of 57 tasks spanning multiple domains such as elementary mathematics, American history, computer science, and law. MMLU-Pro [16], an improved version of MMLU, enhances question quality and accuracy by reducing noise and providing a more detailed assessment of models’ reasoning abilities. MMLU-Pro+ [124] extends its predecessor by evaluating shortcut learning and advanced reasoning capabilities in LLMs. MMLU-Pro+ retains the challenging nature of MMLU-Pro and enhances the assessment of model discernment, especially in situations where multiple correct answers are possible. MMLU-Redux [125] improves the quality and precision of questions through careful curation, leading to a more accurate evaluation.
In contrast, BBH (Big-Bench Hard) is a subset of BIG-Bench, focusing on the most challenging tasks that require multi-step reasoning, spanning a broad spectrum of fields such as mathematics, logic, and commonsense reasoning, aiming to evaluate models’ performance in complex tasks [64]. ARC-C (AI2 Reasoning Challenge - Challenge Set) is dedicated to testing models’ ability to answer complex scientific questions that require logical reasoning, covering science questions from elementary to high school levels, with the goal of assessing models’ scientific reasoning capabilities [126]. TruthfulQA is designed to evaluate the truthfulness of models when answering questions prone to generating false beliefs and biases, using a series of carefully crafted questions to test the reliability and accuracy [69]. Winogrande is a large-scale coreference resolution task that tests models’ ability to handle contextual understanding in sentences through a series of complex questions [110]. HellaSwag evaluates natural language inference by requiring models to complete paragraphs in a way that necessitates understanding complex details, aimed at assessing models’ commonsense reasoning abilities [68]. Besides, RV-Bench [127] evaluates LLMs’ mathematical reasoning by using random variable questions, which require models to understand the underlying problem structure rather than relying on memorized solutions.
While IQ benchmarks have proliferated, significant challenges persist. First, the "memorization vs. reasoning" dilemma complicates assessment—models often succeed through pattern matching rather than genuine understanding. Second, the rapid capability growth of LLMs has rendered many benchmarks obsolete, creating a "red queen" effect where benchmarks quickly become saturated. Third, most IQ assessments remain narrow in scope, failing to capture the full spectrum of human-like reasoning capabilities. Recent studies reveal that even state-of-the-art models struggle with counterfactual reasoning and maintaining consistency across extended dialogues, highlighting gaps in current evaluation methodologies.
3.2 Professional Expertise Evaluation (PQ)
PQ represents the specialized knowledge and skills that an LLM possesses within a particular area. It is akin to the professional acumen that a human expert might have in a specific field. PQ in LLMs is evident in their ability to provide detailed, accurate, and nuanced information within a specialized domain, such as healthcare, financial. Notably, PQ corresponds to capabilities acquired during supervised fine-tuning, where models develop domain-specific expertise through targeted instruction-response learning, forming the operational foundation for specialized LLM applications.
Table 4 shows recent domain-specific evaluation benchmarks, along with additional comparative dimensions such as the scope of tasks, data sources, and unique contributions. This table excludes the introductory descriptions for brevity and focuses on key attributes that facilitate a comparative analysis.
Table 4: 41 typical Professional Quotient (PQ)-Professional Expertise evaluation benchmarks for LLMs.
Domain Name Institution Scope of Tasks Unique Contributions Url
BLURB [128] Mindrank AI Six diverse NLP tasks, thirteen datasets A macro-average score across all tasks link
Seismometer [129] Epic
Using local data and workflows
patient demographics, clinical interventions, and outcomes link
Healthcare Medbench [130] OpenMEDLab Emphasizes scientific rigor and fairness 40,041 questions from medical exams and reports link
GenMedicalEval [131] E
16 majors, 3 training stages, 6 clinical scenarios
Open-ended metrics and automated assessment models link
PsyEval [132] SJTU Six subtasks covering three dimensions Customized benchmark for mental health LLMs link
Fin-Eva [133] Ant Group Wealth management, insurance, investment research Both industrial and academic financial evaluations link
FinEval [134] SUFE-AIFLM-Lab
Multiple-choice QA on finance, economics, accounting
Focuses on high-quality evaluation questions link
Finance OpenFinData [30] Shanghai AI Lab Multi-scenario financial tasks First comprehensive finance evaluation dataset link
FinBen [135] FinAI 35 datasets across 23 financial tasks
Inductive reasoning, quantitative reasoning
link
LAiW [136] Sichuan University 13 fundamental legal NLP tasks Divides legal NLP capabilities into three major abilities link
Legal LawBench [30] Nanjing University
Legal entity recognition, reading comprehension
Real-world tasks, "abstention rate" metric link
LegalBench [137] Stanford University 162 tasks covering six types of legal reasoning Enables interdisciplinary conversations link
LexEval [138] Tsinghua University
Legal cognitive abilities to organize different tasks
Larger legal evaluation dataset, examining the ethical issues
link
SPEC5G [139] Purdue University security-related text classification and summarization 5G protocol analysis automation link
TeleQnA [37] Huawei(Paris) General telecom inquiries
Proficiency in telecom-related questions
link
OpsEval [140] Tsinghua University Wired network ops, 5G, database ops
Focus on AIOps, evaluates proficiency
link
TelBench [141] SK Telecom Math modeling, open-ended QA, code generation Holistic evaluation in telecom link
Telecom TelecomGPT [142] UAE
Telecom Math Modeling, Open QnA and Code Tasks
Holistic evaluation in telecom link
Linguistic [143] Queen’s University Multiple language-centric tasks zero-shot evaluation link
TelcoLM [144] Orange multiple-choice questionnaires
Domain-specific data (800M tokens, 80K instructions)
link
ORAN-Bench-13K [145] GMU multiple-choice questions
Open Radio Access Networks (O-RAN)
link
Open-Telco Benchmarks [146] GSMA
Multiple language-centric tasks
zero-shot evaluation link
FullStackBench [147] ByteDance
Code writing, debugging, code review
Featuring the most recent Stack Overflow QA.
link
StackEval[148] Prosus AI
11 real-world scenarios, 16 languages
Evaluation across diverse&practical coding environments
link
CodeBenchGen [149] Various Institutions Execution-based code generation tasks
Benchmarks scaling with the size and complexity
link
HumanEval [36] University of Washington rigorous testing
Stricter protocol for assessing correctness of generated code
link
APPS [150] University of California Coding challenges from competitive platforms
Checking problems solving of generated code on test cases
link
Coding MBPP [151] Google Research Programming problems sourced from various origins Diverse programming tasks link
ClassEval [152] Tsinghua University Class-level code generation
Manually crafted, object-oriented programming concepts
link
CoderEval [153] Peking University Pragmatic code generation
Proficiency to generate functional code patches for described issues
link
MultiPL-E [154] Princeton University Neural code generation
Benchmarking neural code generation models
link
CodeXGLUE [155] Microsoft Code intelligence
Wide tasks covering: code-code, text-code, code-text and text-text
link
EvoCodeBench [156] Peking University Evolving code generation benchmark
Aligned with real-world code repositories, evolving over time
link
LiveIdeaBench [157] RUC Evaluates scientific creativity and idea generation Single-keyword prompts across 18 domains link
ScienceAgentBench [158] OSU Data-driven scientific discovery 102 tasks from peer-reviewed publications link
Symbolicregression [159] Amazon Symbolic regression for scientific discovery New datasets and evaluation criteria link
Science DiscoveryWorld [160] AIAI Virtual environment for scientific discovery 120 challenge tasks across 8 topics link
ProtocoLLM [161] UT Austin Formulating domain-specific scientific protocols Pseudocode extraction from biology protocols link
SciSafeEval [162] Zhejiang University Safety alignment in scientific tasks Multi-language evaluation with "jailbreak" feature link
SciAssess [163] DP Technology Evaluates proficiency in scientific literature analysis Memorization, comprehension, and analysis link
SciVerse [164] CUHK Evaluating scientific reasoning abilities Covering physics, chemistry, and biology link
3.2.1 Healthcare
The healthcare domain has seen the development of specialized benchmarks to evaluate LLMs (LLMs) in medical applications, each with unique features contributing to comprehensive evaluation. Seismometer [129] supports continuous monitoring of model performance within local data and workflows, ensuring models remain effective over time. BLURB [128] offers a suite for biomedical NLP tasks using 13 publicly available datasets across 6 diverse tasks. Medbench [30], provides a robust medical LLM evaluation system through 40,041 questions from authentic examination exercises. GenMedicalEval [131] covers 16 major departments with over 100,000 real-world medical cases, while PsyEval [132] is tailored specifically for mental health applications. MedS-Bench [165] introduces a large-scale instruction-tuning dataset MedS-Ins for medicine, comprising 58 medically oriented language corpora, totaling 5M instances with 19K instructions, across 122 tasks, and launches a dynamic leaderboard for MedS-Bench.
3.2.2 Financial
Fin-Eva [133], OpenFinData [30], and FinEval [134] Finben [135] provide structured evaluations of LLMs’ financial capabilities. Fin-Eva evaluates LLMs using over 13,000 multiple-choice questions covering various financial scenarios. OpenFinData includes diverse data types from business scenarios, ensuring practical applicability. FinEval focuses on high-quality multiple-choice questions that adhere to professional standards. Practical guidance may emphasize selecting benchmarks that not only cover a broad range of scenarios but also integrate into existing financial operations.
3.2.3 Legal
Benchmarks like LAiW [136], LawBench [30], and LegalBench [137] offer detailed assessments in legal contexts. LAiW divides legal NLP into three categories, including complex legal application tasks. LawBench simulates judicial cognition through twenty tasks and introduces an "abstention rate" metric [166]. LegalBench [166] encompasses 162 tasks covering 6 types of legal reasoning. These benchmarks collectively aim to bridge the gap between legal professionals and LLM developers, promoting transparency and rigor in evaluations. The introduction of metrics like the “abstention rate" in LawBench [30] adds a layer of nuance to evaluating LLMs’ ability to handle ambiguous or complex instructions.
3.2.4 Telecommunications
The benchmarks such as TeleQnA [37], TelBench [141], and TelecomGPT [142] address unique challenges in evaluating LLMs. TeleQnA [37] evaluates LLMs using 10,000 telecom-related Q&A pairs. TelBench [141] extends existing benchmarks with new tasks like Telecom Math Modeling and Code Tasks. TelecomGPT [142] proposes adaptation pipelines for general-purpose LLMs to telecom-specific models. Besides, interdisciplinary OpsEval [140] evaluates LLMs in wired network operations, 5G, and database operations, supporting evaluations in English and Chinese.
3.2.5 Coding
The evaluation within the coding domain is a critical area that has obtained significant attention due to its potential impact on software development practices and automated programming tools [36]. The benchmarks designed for this purpose aim not only to assess the syntactic correctness of generated code, but also to evaluate more complex aspects such as semantic accuracy, functionality, and efficiency. We highlight several key points regarding the current state and future directions of LLM evaluation for coding.
Existing benchmarks cover a spectrum of tasks, from syntactic correctness to semantic accuracy, functionality, and efficiency. FullStackBench [147] offers comprehensive real-world scenarios across multiple programming languages, while CodeBenchGen [149] focuses on execution-based code generation tasks and scales with the complexity of programming challenges. EvoCodeBench [167] evolves over time to reflect contemporary coding practices, and HumanEval [36] provides a strict evaluation protocol for code correctness. APPS [150] assesses algorithmic problem-solving skills, and MBPP [151] evaluates basic programming tasks. CoderEval [153] emphasizes generating functional code patches, MultiPL-E [154] offers a scalable framework for neural code generation, and CodeXGLUE [155] covers a range of code intelligence tasks.
Specifically, FullStackBench [147] and CodeBenchGen [149] offer coverages of coding environments, but their static nature may limit their ability to adapt to evolving coding standards. EvoCodeBench [167] addresses this by evolving over time, ensuring that benchmarks remain relevant to contemporary practices. HumanEval [36] and APPS [150] focus on code correctness and efficiency, making them essential for verifying practical utility. MBPP [151] evaluates basic programming skills, while CoderEval [153], MultiPL-E [154], and CodeXGLUE [155] address specific aspects like functional code patches, neural code generation, and code intelligence.
3.2.6 Software
In software engineering, benchmarks like SWE-bench [168], Owl-Bench [169] and CodeMMLU [170] provide structured assessing approaches in software development. SWE-bench [168] evaluates LLMs’ ability to resolve real-world GitHub issues, while Owl-Bench assesses their proficiency in software documentation [169]. CodeMMLU [170] includes 10K questions sourced from diverse domains, encompassing tasks like code analysis, defect detection, and software engineering principles across programming languages.
These benchmarks collectively cover a broad spectrum of software engineering tasks, from operations management to issue resolution and documentation. The comparison highlights the importance of task-oriented evaluations and practical application scenarios, ensuring that LLMs can effectively assist in real-world software development processes.
Table 5: 37 typical Emotional Quotient (EQ)-Alignment Ability evaluation benchmarks for LLMs (zoom in).
Name Year Task Type Institution Category Datasets Url
DiffAware [171] 2025 Bias Stanford General Bias 8 datasets link
CASE-Bench [172] 2025 Safety Cambridge Context-Aware Safety CASE-Bench link
Fairness [173] 2025 Fairness PSU Distributive Fairness - -
HarmBench [174] 2024 Safety UIUC Adversarial Behaviors 510 link
SimpleQA [175] 2024 Safety OpenAI Factuality 4,326 link
AgentHarm [176] 2024 Safety BEIS Malicious Agent Tasks 110 link
StrongReject [177] 2024 Safety dsbowen Attack Resistance n/a link
LLMBar [178] 2024 Instruction Princeton Instruction Following 419 Instances link
AIR-Bench [179] 2024 Safety Stanford Regulatory Alignment 5,694 link
TrustLLM [180] 2024 General TrustLLM Trustworthiness 30+ link
RewardBench [29] 2024 Alignment AIAI Human preference RewardBench link
EQ-Bench [181] 2024 Emotion Paech Emotional intelligence 171 Questions link
Forbidden [182] 2023 Safety CISPA Jailbreak Detection 15,140 link
MaliciousInstruct [183] 2023 Safety Princeton Malicious Intentions 100 link
SycophancyEval [184] 2023 Safety Anthropic Opinion Alignment n/a link
DecodingTrust [185] 2023 Safety UIUC Trustworthiness 243,877 link
AdvBench [186] 2023 Safety CMU Adversarial Attacks 1,000 link
XSTest [187] 2023 Safety Bocconi Safety Overreach 450 link
OpinionQA [188] 2023 Safety tatsu-lab Demographic Alignment 1,498 link
SafetyBench [189] 2023 Safety THU Content Safety 11,435 link
HarmfulQA [190] 2023 Safety declare-lab Harmful Topics 1,960 link
QHarm [174] 2023 Safety vinid Safety Sampling 100 link
BeaverTails [191] 2023 Safety PKU Red Teaming 334,000 link
DoNotAnswer [192] 2023 Safety Libr-AI Safety Mechanisms 939 link
AlignBench [25] 2023 Alignment THUDM Alignment, Reliability Various link
IFEval [24] 2023 Instruction Google Instruction Following 500 Prompts link
ToxiGen [193] 2022 Safety Microsoft Toxicity Detection 274,000 link
HHH [194] 2022 Safety Anthropic Human Preferences 44,849 link
RedTeam [195] 2022 Safety Anthropic Red Teaming 38,961 link
BOLD [196] 2021 Bias Amazon Bias in Generation 23,679 link
BBQ [197] 2021 Bias NYU Social Bias 58,492 link
StereoSet [198] 2020 Bias McGill Stereotype Detection 4,229 link
ETHICS [199] 2020 Ethics Berkeley Moral Judgement 134,400 link
ToxicityPrompt [200] 2020 Safety AllenAI Toxicity Assessment 99,442 link
CrowS-Pairs [201] 2020 Bias NYU Stereotype Measurement 1,508 link
SEAT [202] 2019 Bias Princeton Encoder Bias n/a link
WinoGender [203] 2018 Bias UMass Gender Bias 720 link
3.2.7 Science
It is a critical area where LLMs have the potential to significantly impact research and discovery processes [157]. Evaluating LLMs in this domain requires specialized benchmarks that assess their ability to understand, generate, and apply scientific knowledge across diverse fields such as biology, chemistry, physics, and medicine. This section provides an overview of prominent evaluation benchmarks designed to assess LLMs’ capabilities in scientific tasks.
Key benchmarks—LiveIdeaBench [157], ScienceAgentBench [158], Symbolicregression [159], DiscoveryWorld [160], ProtocoLLM [161], and SciSafeEval [162]—are pivotal for LLMs in scientific domains. LiveIdeaBench [157] assesses models’ scientific creativity and divergent thinking across four dimensions (originality, feasibility, fluency, flexibility) using single-keyword prompts. SciAssess [163] evaluates LLMs’ proficiency in scientific literature analysis, including memorization and comprehension tasks. SciVerse [164], a multi-modal benchmark, tests scientific reasoning abilities with annotated Q&A samples. DiscoveryWorld [160] benchmarks agents’ ability to perform novel scientific discovery cycles. ProtocoLLM [161] evaluates the ability to formulate domain-specific scientific protocols. SciSafeEval [162] ensures safety alignment across scientific tasks, introducing a “jailbreak" feature to test defenses against malicious intentions.
Collectively, these benchmarks provide a comprehensive framework for evaluating LLMs’ capabilities in the science domain. They highlight not only the importance of scientific creativity and literature analysis but also emphasize practical aspects such as hands-on experimentation, hypothesis testing, and ethical considerations. For instance, LiveIdeaBench [157] and SciAssess [163] offer unique methodologies for assessing divergent thinking and innovative idea generation, indicating that LLMs require distinct evaluation approaches beyond traditional memory and understanding. On the other hand, DiscoveryWorld [160] and ProtocoLLM [161] focus on practical skills, underscoring the significance of experimental design and hypothesis formation, which are essential for cultivating LLMs’ actual research capabilities. Furthermore, SciVerse [164] and SciSafeEval [162] extend the evaluation scope to include multi-modal reasoning and safety alignment, ensuring that LLMs can effectively handle complex datasets while adhering to ethical standards. Collectively, these benchmarks guide the development of more advanced LLMs, ultimately contributing to accelerating scientific innovation and discovery.
3.3 Alignment Ability Evaluation (EQ)
The concept of Alignment Ability, often referred to as Emotional Quotient (EQ) in the context of LLMs, is a critical aspect of evaluating how well these models can understand and appropriately respond to the emotional and social nuances within human interactions. This evaluation is essential for ensuring that LLMs not only generate text that is coherent and relevant but also that they do so in a manner that is empathetic, culturally sensitive, and ethically sound [204]. Specifically, EQ corresponds to capabilities refined through reinforcement learning from human feedback, where models learn to align outputs with human values, ensuring socially appropriate and ethically sound interactions.
As shown in Table 5, benchmarks have been developed to assess the EQ of LLMs, each focusing on different aspects of emotional intelligence. For instance, EQ-Bench [181] is a notable benchmark specifically designed to evaluate the emotional intelligence of LLMs. It challenges the models to predict the intensity of emotional states of characters in a dialogue, thereby assessing their ability to understand complex emotions and social interactions. The EQ-Bench dataset consists of 171 carefully crafted questions, providing a robust framework for measuring the emotional acumen of LLMs. Meanwhile, AlignBench includes a comprehensive multi-dimensional approach to evaluating the alignment of LLMs with human intent [25], it encompasses a wide range of categories, including reliability, and it uses a combination of 683 real-scenario rooted queries and corresponding human-verified references to ensure that the evaluation reflects actual usage contexts. This benchmark allows for a nuanced assessment of model performance across various dimensions, such as creativity, logic, and sensitivity.
RewardBench [29] and TrustLLM [180] are also noteworthy, as they focus on different facets of alignment. RewardBench evaluates the reward modeling capabilities of LLMs, which is crucial for understanding and following instructions, while TrustLLM measures the trustworthiness of models, an essential component of user confidence and safety. These benchmarks, along with others like IFEval [24] and LLMBar [178], which concentrate on instruction following, provide a comprehensive suite of tools for researchers and developers to measure and improve the alignment of LLMs with human expectations. Besides, the Fairness benchmark [173] and CASE-Bench [172] both highlight the importance of aligning LLMs with human values. The Fairness benchmark evaluates LLMs’ alignment with distributive fairness concepts like equitability and envy-freeness, revealing a lack of alignment with human preferences. CASE-Bench focuses on safety, integrating context into safety assessments and showing context’s significant influence on human judgments. Both underscore the need for LLMs to better align with societal norms [205].
4 Value-oriented Evaluation of LLMs
Extant works predominantly employ conventional performance metrics to assess LLMs. However, these metrics are frequently insufficient to encapsulate the complex societal, economic, ethical, and environmental repercussions of deploying LLMs. Recent studies have begun to explore alternative evaluation frameworks that consider a broader spectrum of impacts, signaling a shift towards more holistic assessments. As shown in Fig. 4, this section delves into a value-oriented evaluation framework for LLMs, which transcends conventional performance benchmarks to encompass a holistic assessment including economic, social, ethical, and environmental considerations. By advocating for an evaluation approach that not only quantifies technical proficiency but also qualifies the broader implications of LLM deployment, this paper aims to contribute to the discourse on responsible AI development.
Figure 4: Value-oriented Evaluation for LLMs.
Economic Value
We give some key metrics: Cost-Benefit Ratio (CBR): This metric evaluates the ratio of the benefits derived from the model to the costs incurred in its development and deployment. A higher CBR indicates a more economically viable solution. Return on Investment (ROI): it measures the financial return generated by the model relative to the initial investment. It provides a clear indication of the model’s profitability and long-term financial viability. Productivity Improvement (PI): PI assesses the extent to which the model enhances productivity in specific application domains. For instance, in a business setting, an LLM that automates customer service can significantly reduce response times and improve efficiency. Market Acceptance (MA): Market acceptance is a qualitative metric that gauges the level of adoption and user satisfaction with the model. High market acceptance suggests that the model meets the needs and expectations of its target audience.
Social Value
The following metrics are used to evaluate social value: User Satisfaction (US): User satisfaction is a direct measure of how well the model meets the needs and preferences of its users. Surveys and feedback mechanisms can be employed to gather this data. Knowledge Dissemination Efficiency (KDE): it measures the effectiveness of the model in spreading knowledge and information. In educational settings, for example, an LLM that can generate high-quality learning materials can significantly enhance the dissemination of knowledge. Public Service Improvement (PSI): it evaluates the extent to which the model improves the quality and efficiency of public services. Case studies and expert reviews can provide insights into the model’s impact on public service delivery. Education Quality Improvement (EQI): it assesses the contribution of the model to enhancing the quality of education. Metrics such as student performance and teacher feedback can be used to quantify this improvement.
Ethical Value
Ethical considerations are paramount in the deployment of LLMs, as these models can have significant implications for fairness, transparency, and privacy. The following metrics are used to evaluate ethical value: Fairness (F): Fairness ensures that the model performs equitably across different demographic groups. Statistical tests and bias detection methods can be used to identify and mitigate any disparities. Transparency (T): Transparency refers to the model’s ability to provide understandable and clear explanations for its decisions. Expert reviews and user comprehension tests can help assess the model’s transparency. Privacy Protection (PP): Privacy protection measures the model’s capability to safeguard personal data. Security audits and compliance checks are essential for ensuring that the model adheres to privacy regulations. Bias Detection (BD): Bias detection involves identifying and quantifying any biases present in the model. Regular audits and bias mitigation strategies are necessary to maintain the model’s ethical integrity.
Environmental Value
It considers the ecological impact of LLMs, including energy consumption and carbon footprint. The following metrics are used: Energy Efficiency (EE): EE measures the energy consumption of the model during operation. Carbon Footprint (CF): CF quantifies the total carbon emissions associated with the model’s lifecycle, from development to deployment. Reducing the carbon footprint is crucial for mitigating the environmental impact of AI technologies. Sustainability (S): Sustainability evaluates the long-term environmental and social impact of the model. Life cycle assessments and future projections can provide a comprehensive view of the model’s sustainability.
5 LLM System or Application Evaluation
In this section, we delve into the intricacies of evaluating LLM systems and applications, exploring the methodologies, metrics, and benchmarks that are pivotal in ensuring the advancement and responsible deployment of these powerful AI tools. It also focuses on three pivotal areas: Retrieval-Augmented Generation (RAG), AI Agents, and Chatbots.
Table 6: Comparison of Retrieval-Augmented Generation (RAG) Evaluation Frameworks.
Name Institute Feature Domain Evaluation Criteria Url
RAGAS [206] Exploding Gradients Automated Evaluation QA Answer Relevance, Context Relevance, Faithfulness code
BER [207] NAVER Benchmarking RAG QA Consistency in benchmarking RAG pipelines code
CRAG [208] Meta Reality Labs Factual QA Benchmark QA Diverse questions across multiple domains code
rag-llm-hub [209] RAGA-AI Comprehensive Evaluation Toolkit Various Multiple aspects including relevance, quality, safety code
ARES [210] Stanford Automatic Evaluation for RAG QA Context Relevance, Answer Faithfulness code
RGB [211] CAS Performance, Robustness QA Counterfactual Robustness, Information Integration code
BEIR [212] UKP-TUDA Out-of-distribution, Zero-shot QA, Bio-Medical IR Out-of-distribution, zero-shot code
ALCE [213] Princeton NLP Citation, Hallucination Generate with Citations Citation Quality, Correctness, Fluency code
KITAB [214] Microsoft Constraint IR Constraint IR All correct, Completeness, etc. code
NoMIRACL [215] Project MIRACL Multilingual Robustness Evaluation Error Rate, Hallucination Rate code
CRUD-RAG [216] IAAR-Shanghai CRUD Operations QA, Hallucination Creative Generation, Error Correction, etc. code
Table 7: Main Evaluation Metrics for Assessing RAG.
Metrics Details Reference
Faithfulness Assesses the factual alignment between the generated response and the provided context. Link
Answer Relevance Examines the degree to which the generated response is relevant to the given prompt. Link
Context Precision Determines if all context items relevant to the ground truth are appropriately ranked. Link
Context Relevancy Evaluates the relevance of the retrieved context based on the question and contexts. Link
Context Recall Assesses how well the retrieved context matches the annotated answer, considered as the ground truth. Link
Answer Semantic Similarity Measures the semantic closeness between the generated answer and the ground truth. Link
Answer Correctness Evaluates the accuracy of the generated answer in comparison to the ground truth. Link
5.1 RAG Evaluation
Retrieval-Augmented Generation (RAG) has emerged as a pivotal approach to enhancing the capabilities of LLMs by integrating retrieval mechanisms with generative processes [28]. The evaluation of RAG models focuses on the model’s ability to incorporate retrieved information seamlessly into its responses[206]. This assessment goes beyond merely judging the quality of the generated text, it also scrutinizes the precision and relevance of the retrieved data, alongside how well this information complements and enriches the final output. Key performance indicators for RAG systems typically encompass the retrieval process’s accuracy and completeness, as well as the logical consistency and contextual appropriateness of the augmented content. Table 6 provides a comprehensive overview of various benchmarks designed to assess the performance of RAG systems across diverse domains.
The diversity of evaluation aspects and metrics employed by these frameworks highlights the multifaceted nature of RAG assessment. For instance, RAGAS from Exploding Gradients focuses on automated evaluation through customized metrics that measure answer relevance, context relevance, and faithfulness [206]. It is particularly valuable for its ability to evaluate the alignment between retrieved contexts and generated answers, ensuring that the output remains grounded in factual information. Similarly, BERGEN emphasizes consistency in benchmarking RAG pipelines, addressing the challenge of inconsistent evaluations that can hinder comparative analysis [217]. By leveraging HuggingFace for reproducibility and integration, BERGEN facilitates a standardized approach to evaluating RAG systems, thereby promoting transparency and comparability in research findings. Table 7 encapsulates a range of evaluation metrics essential for assessing the performance of LLMs (LLMs). Each metric serves a distinct purpose, contributing to a comprehensive evaluation framework that ensures models are not only technically proficient but also contextually relevant and factually accurate. These metrics together form a robust evaluation framework that supports the development and deployment of LLMs by offering detailed insights into their performance across dimensions.
On the other hand, CRAG introduces a benchmark to simulate web and Knowledge Graph (KG) search, covering a wide array of question types and domains [208]. Such extensive coverage allows researchers to explore the robustness and versatility of RAG systems under varying conditions. In contrast, raga-llm-hub offers a comprehensive toolkit with over 100 evaluation metrics, focusing on multiple dimensions such as relevance, quality, safety, and more [218]. This breadth of assessment ensures that developers can thoroughly evaluate LLMs and RAG applications, identifying areas for improvement and optimizing performance.
For practical use, ARES exemplifies this transition by providing an automatic evaluation framework that includes human-annotated datasets for scoring context relevance, answer faithfulness, and answer relevance [210]. The use of annotated data enhances the reliability of evaluations, offering insights into both the strengths and weaknesses of RAG systems. Moreover, RGB [211] focuses on four fundamental capabilities: negative rejection, noise robustness, counterfactual robustness, and information integration. BEIR focus on out-of-distribution and zero-shot tasks underscores the importance of adaptability in RAG systems, preparing them for scenarios where prior knowledge may be limited [212]. Meanwhile, ALCE [213] emphasizes on citation quality and correctness addresses concerns about hallucinations, ensuring that generated content adheres to established facts and sources.
Table 8: Comprehensive Comparison of Agent Evaluation Benchmarks.
Name Institutions Domain Metrics Tool Interaction Multi-Agent Role-Playing
SuperCLUE-Agent[219] CLUE Various Chinese tasks Core abilities, 10 fundamental tasks Limited No No
AgentBench[220] THU Coding, Gaming, Web Success rates, F1 scores Yes No No
API-Bank[221] Alibaba Tool invocation scenarios API search accuracy, response quality Yes No No
AgentBoard[222] UHK Multi-task Process rate, grounding accuracy, sub-capabilities Yes Yes No
MetaTool[223] Lehigh University Tool invocation Similar tool choice, context-specific, reliability, multi-tool Yes No No
Agents That Matter[224] Princeton N/A Cost-effectiveness, joint optimization No No No
PersonaGym[225] CMU Role-playing scenarios PersonaScore No No Yes
MMRole[226] RUC Multimodal role-playing Instruction Adherence, Fluency, Coherency, Consistency No No Yes
GLEE [227] IIT Economic contexts Parameterization, degrees of freedom Yes Yes Yes
BFCL [228] UC Berkeley Function-calling tasks Success rate in function calls, parallel execution Yes No No
ToolLLM [65] OpenBMB Real-world APIs Instruction tuning effectiveness Yes No No
ToolBench [229] SambaNova Systems Tools for real-world tasks Tool manipulation capability Yes No No
Webarena [230] WebArena-X Web-based environments Task completion on the web Yes No No
5.2 Agent Evaluation
The advent of LLMs has led to advancements in AI Agents capable of autonomously interacting with various environments and tools. To ensure that these agents meet the desired standards, a variety of evaluation frameworks have emerged [220, 221, 219, 222]. Each framework targets different aspects of Agent performance, such as tool usage, decision-making, role-playing, and multi-modal interaction. Table 8 compares several key benchmarks across multiple dimensions, highlighting their unique contributions to the field.
AgentBench [220] and API-Bank [221] emphasize evaluating Agents across diverse real-world scenarios, including coding, gaming, web interactions, and tool invocations. This broad scope ensures that Agents are tested under conditions closely resembling their intended operational environments, providing valuable feedback on their generalization capabilities.
Metrics play a crucial role in assessing Agent performance. For example, AgentBoard [222] introduces novel metrics such as process rate and grounding accuracy, offering deeper insights into how effectively Agents handle complex tasks. Meanwhile, MMRole [226] evaluates multimodal interaction through detailed criteria considering both textual and visual elements, ensuring a more holistic assessment.
Moreover, new entries like BFCL [228] focus on function-calling tasks, including multi-task and parallel function calls, challenging the Agents’ ability to handle complex logic. ToolLLM [65] enables LLMs to master over 16,000 real-world APIs, while ToolBench [231] assesses the capability of Agents to manipulate software tools used in real-world tasks. Webarena [230] creates realistic web environments for Agents to complete various web-based tasks. The GLEE [227] framework focuses on agents’ behavior within economic contexts, using parameters such as parameterization, degrees of freedom, and economic measures to evaluate agent performance. This highlights the importance of understanding societal and economic activities. The lack of standardized evaluation methods remains a challenge. Frameworks like PersonaGym [225] introduce scoring systems, such as PersonaScore, which could pave the way for establishing industry-wide standards.
Table 9: Comprehensive Evaluation of LLM-based Chatbot Frameworks.
Name Feature Domain Evaluation Criteria Metric
ChatBotBenchmark [232] Multi-turn chatting capability Dialogue systems Consistency, Coherence BLEU, ROUGE
MT-Bench-101 [233] Fine-grained abilities Dialogue systems Turn-taking skills, Error handling Accuracy, F1 score
FairMT-Bench [234] Fairness in conversations Dialogue systems Bias detection, Fairness Fairness index, Bias rate
MT-Eval [235] Interaction patterns Human-LLM interactions Interaction quality, Error propagation Interaction score, Error rate
MINT [236] Problem-solving capabilities Multi-turn interactions Tool usage, Feedback integration Success rate, Efficiency
Chatbot Arena [66] Competitive LLM comparison platform Dialogue systems Human preference Preference scores
MixEval [237] Dynamic benchmark from mixtures Multi-turn dialogues Crowd wisdom Derived metrics
WildChat [238] 1M real-world ChatGPT interactions Dialogue systems User behavior Usage patterns
MT-Bench [33] Multi-turn follow-up questions Dialogue systems Dialogue quality Human judgments
CoQA [239] Multi-turn QA Question answering Answer coherence F1 score, BLEU
QuAC [240] Contextual student-teacher QA Question answering Contextual understanding F1 score, ROUGE
5.3 ChatBot Evaluation
The assessment of modern chatbot systems, particularly those based on LLMs, requires multidimensional frameworks addressing linguistic coherence, contextual understanding, and ethical considerations (Table 9). As conversational AI evolves from single-turn responses to multi-party dialogues, traditional evaluation metrics such as BLEU [39] and ROUGE [41] prove insufficient for capturing the complexity of human-like interactions. This section analyzes state-of-the-art benchmarks across 3 critical dimensions: dialogue quality, fairness and human interaction patterns.
Dialogue Quality Assessment: it focuses on structural, linguistic, and contextual dimensions. BotChatBenchmark [232] introduces the ChatSEED methodology, where real-world dialogue snippets serve as prompts for LLMs to generate full-length conversations. Using GPT-4 as a meta-judge, this framework reveals significant performance disparities: while GPT-4 achieves top consistency with human dialogues, open-source models like Llama2-70B exhibit suboptimal verbosity errors. MT-Bench-101 [233] extends this analysis through a three-tier taxonomy covering 13 tasks, exposing critical failure modes in error recovery and instruction-following. Besides, the MT-Bench framework [33] establishes human judgment standards, demonstrating that crowd-sourced evaluations correlate with expert assessments. For question-answering systems, CoQA [239] and QuAC [240] employ F1/ROUGE metrics, revealing that models struggle with pronoun resolution.
Fairness Evaluation: FairMT-Bench [234] constructs a 10K-dialogue dataset spanning gender, ethnicity, and occupational biases, showing that LLMs exhibit up to 37% performance variance across sensitive scenarios. MixEval [237] addresses dataset bias through a meta-benchmarking approach, aggregating samples from existing benchmarks to create dynamic criteria. Their “wisdom of crowds" metric reveals that model rankings change over benchmark mixtures.
Human Interaction Patterns: Human interaction analysis emphasizes real-world dynamics. Chatbot Arena [66] collects 33K competitive dialogues through a crowdsourced platform, demonstrating that closed-source models (e.g., GPT-4) outperform open-source alternatives in user preference scores. MT-Eval [235] identifies four interaction patterns—recollection, expansion, refinement, and follow-up—showing that error propagation increases in multi-turn settings. WildChat [238] provides unprecedented scale with 1M ChatGPT interactions, revealing different user behaviors.
6 Challenges and Outlook
We propose a six-tiered challenges and future opportunities: starting from foundational methodological concerns (statistical rigor and reproducibility), advancing through technical evaluation complexities (composite metrics and interpretability), extending to application-level considerations (user experience and human-in-the-loop assessment), encompassing system-level evaluation (pragmatic system analysis and failure exploration), adapting to evolutionary dynamics (dynamic evaluation mechanisms), and ultimately reaching value-oriented dimensions (economic, social, ethical, and environmental impacts). This structure reflects how LLM evaluation must evolve from purely technical assessments toward holistic frameworks.
Enhanced Statistical Analysis for LLM Evaluation
Current evaluation practices suffer from a critical methodological gap: the lack of rigorous statistical foundations necessary for reliable performance assessment. Most benchmarks report point estimates without confidence intervals, making it difficult to determine whether observed performance differences represent genuine capability improvements or merely statistical noise. Integrating rigorous statistical methods is essential to transform LLM evaluation from simplistic scoring to scientifically valid methodology for reliable model development.
Composite Evaluation/Ranking Systems
Developing composite and comprehensive evaluation/ranking systems represents the necessary evolution beyond basic statistical rigor. Current evaluation methods often focus on specific tasks or benchmarks, which may not fully capture the multifaceted capabilities of LLMs. A composite system that integrates various metrics and evaluation criteria can provide a more nuanced and comprehensive assessment.
Interpretability and Explainability
One fundamental challenge in evaluating LLMs is the alignment between the fine-grained decision-making logic of the models and human cognition. Current evaluation practices often focus on the correctness of the output, merely addressing hallucination and value alignment issues. However, in practical industrial applications, the crux of assessing the credibility of LLMs lies in the correctness of the underlying decision logic that leads to the output. This is particularly challenging because, even though LLMs may exhibit high accuracy on specific tasks, their internal decision logic can be highly chaotic and misaligned with human reasoning. Developing explainable AI (XAI) techniques specifically tailored for LLMs can enhance transparency and facilitate better human-AI collaboration.
User-Centric Experience as a Benchmark
Moving beyond purely technical assessments, user-centric experience represents a crucial application-level consideration. Traditional benchmarks often focus on technical performance metrics, which may not fully capture the user’s perspective. Incorporating user feedback and usability testing can provide more valuable insights into the practical utility and user satisfaction of LLMs. This can be achieved via user studies, surveys, and interactive sessions with qualitative data on user experiences.
Human in the Loop Evaluation (HITL)
Human in the Loop Evaluation extends user-centric assessment into a more sophisticated system-level framework. This approach is crucial for addressing the limitations of automated evaluation methods. HITL involves human evaluators who can provide subjective judgments and context-specific insights that automated systems may miss. HITL enhances the relevance and reliability of evaluations, ensuring that models are judged based on their actual utility rather than just theoretical benchmarks. Furthermore, the Arena Module concept addresses limitations inherent in static leaderboards by offering ongoing assessments that evolve with user interaction, providing a dynamic and realistic evaluation environment in actual usage contexts.
Analytical Failure Exploration
Understanding the root causes of failures represents a deeper layer of system evaluation that moves beyond surface-level performance metrics. Analytical failure exploration involves identifying and analyzing the specific reasons why an LLM fails in certain tasks. This can be achieved through techniques such as error analysis, case studies, and post-hoc explanations. By pinpointing the underlying issues, researchers can develop targeted interventions to address these weaknesses. Additionally, sharing failure cases and their analyses can foster a collaborative environment where the community can learn from each other’s experiences and collectively improve LLMs. This approach moves evaluation from merely identifying what fails to understanding why it fails, enabling more meaningful improvements in model design and deployment strategies.
Dynamic Evaluation
It represents a critical shift from one-time assessment to continuous evaluation. Dynamic evaluation ensures that LLMs are assessed under realistic and up-to-date conditions, promoting continuous improvement and innovation.
Superior Value-Oriented Evaluation
The highest tier of evaluation considerations would encompasses value-oriented dimensions that transcend technical performance to consider broader societal implications. Implementing a value-oriented evaluation framework requires a multi-faceted implementation, combining quantitative and qualitative analysis, data collection, expert reviews, and user feedback. This represents the natural culmination of evaluation considerations, from technical assessment to societal impact.
7 Conclusions
This survey repositions LLM evaluation beyond benchmark-centric approaches by introducing an anthropomorphic framework that bridges the critical gap between technical performance and real-world impact. We pioneer a holistic IQ-EQ-PQ-VQ taxonomy—integrating General Intelligence, Alignment Ability, Professional Expertise, and Value Quotient, that transcends fragmented metrics to capture what LLMs know, how they apply knowledge, why their outputs resonate with human values, and how they contribute to societal well-being. Critically, this taxonomy reflects the developmental trajectory of LLMs themselves, with IQ corresponding to pre-training knowledge acquisition, PQ emerging from supervised fine-tuning, and EQ cultivated through reinforcement learning—providing not just an evaluation framework but a diagnostic lens for model development. The systematic analysis of over 200 benchmarks across six dimensions that reveals hidden interconnections and critical gaps, we present a modular evaluation architecture with six interconnected components that provides practitioners with actionable guidance for end-to-end evaluation pipelines.
Appendix A Appendix
The following appendix provides supplementary information regarding the evaluation of LLMs and highlights some of the most prominent LLMs currently available. It aims to offer a comprehensive overview of the methodologies used to assess these models and to showcase examples of leading models in the field.
A.1 Evaluation Methodology
A.1.1 Metric-centered Evaluation
It focuses on quantifying the performance of LLMs (LLMs) using standardized metrics. Common metrics include BLEU, ROUGE, METEOR, and BERTScore, each capturing different aspects of text quality and relevance. For example, BLEU measures the precision of n-grams in generated text compared to reference texts, while ROUGE focuses on recall, assessing how well the generated text captures key ideas from the reference. BERTScore, on the other hand, leverages contextual embeddings to evaluate semantic similarity, providing a more nuanced assessment of text quality. These metrics are essential for benchmarking and comparing LLMs across various tasks and datasets.
A.1.2 Human-centered Evaluation
Human-centered evaluation involves human judges assessing the quality, relevance, and naturalness of LLM-generated text. This approach complements metric-centered evaluation by capturing subjective aspects that automated metrics may miss. For example, humans can evaluate whether generated text is coherent, contextually relevant, and free from biases. Human evaluation can also involve tasks such as rating the faithfulness of generated text to the input context or assessing the overall quality of generated summaries. This method is particularly important for evaluating the practical utility of LLMs in real-world applications.
A.1.3 Model-Centric Evaluation
It focuses on the internal mechanisms and capabilities of LLMs. This includes analyzing the model’s architecture, training process, and the quality of its embeddings. For example, evaluating the alignment between the model’s decision logic and human reasoning is crucial for ensuring that LLMs produce outputs that are not only correct but also interpretable. Techniques e.g. feature importance analysis and attention mechanisms can provide insights into the model’s decision-making process, helping to identify potential biases or areas for improvement.
A.2 Prominent LLMs
Several prominent LLMs have emerged in recent years, each with unique capabilities and applications. For example, GPT-4 from OpenAI has demonstrated advanced capabilities in natural language understanding and generation. Other notable models include Meta’s Llama series and Alibaba’s Qwen series, which have been fine-tuned for various NLP tasks. These models are evaluated using a combination of intrinsic metrics (such as, perplexity, accuracy) and extrinsic metrics (such as, performance on specific tasks) to assess their overall effectiveness. The choice of LLM often depends on the specific application, with each model offering trade-offs in terms of performance, computational efficiency, and ease of use. Table 10 demonstrates list of prominent LLMs (published after 2022 and model parameters over 1B) and their basic information.
Table 10: List of Prominent LLMs and their basic information ( accurate as of August 20, 2025), includes representative models and does not encompass all available models.
Model Date Organization Country Para (B) Arena Elo
Qwen3-235B-A22B 2025-07-22 Alibaba China 235 1422
Grok-4 2025-07-09 xAI USA - 1425
Gemini 2.5 Pro 2025-06-05 Google USA - 1457
DeepSeek-R1 2025-05-28 DeepSeek China 671 1417
Claude Sonnet 4 2025-05-22 Anthropic USA - -
o3 2025-04-16 OpenAI USA - 1445
Llama 4 Maverick 2025-04-08 Meta AI USA - -
Llama 3.1 Nemotron 2025-04-07 NVIDIA USA 253 1345
GPT-4.5 2025-02-27 OpenAI USA - 1439
DeepSeek-V3 2024-12-24 DeepSeek China 671 1317
Llama 3.3 2024-12-06 Meta AI USA 70 1274
Hunyuan-Large 2024-11-06 Tencent China 389 1250
Doubao-pro 2024-10-28 ByteDance China - -
Palmyra X 004 2024-10-09 Writer USA - -
Qwen2.5-72B 2024-09-19 Alibaba China 73 1283
Jamba 1.5-Large 2024-08-22 AI21 Labs Israel 398 1305
AFM-on-device 2024-07-29 Apple USA - -
Mistral Large 2 2024-07-24 Mistral AI France 123 1276
Llama 3.1-405B 2024-07-23 Meta AI USA 405 1269
DeepSeek-Coder-V2 2024-06-17 DeepSeek China 236 1214
Nemotron-4 340B 2024-06-14 NVIDIA USA 340 1209
Qwen2-72B 2024-06-07 Alibaba China 73 1187
Llama 3-70B 2024-04-18 Meta AI USA 70 1248
ReALM 2024-03-29 Apple USA - -
DBRX 2024-03-27 Databricks USA 132 1103
AraMCO 2024-03-04 Saudi Aramco SA 250 -
MegaScale 2024-02-23 ByteDance China 530 -
Aya 2024-02-12 Cohere Multi 13 1179
Qwen1.5-72B 2024-02-04 Alibaba China 72 1118
Palmyra X 003 2024-01-01 Writer USA 72 -
Mixtral 8x7B 2023-12-11 Mistral AI France 467 1148
Llama Guard 2023-12-07 Meta AI USA 70 1206
Qwen-72B 2023-11-30 Alibaba China 72 1187
PPLX-70B 2023-11-29 Perplexity USA 70 1081
Nemotron-3-8B 2023-11-15 NVIDIA USA 8 -
Grok-1 2023-11-04 xAI USA 314 1266
BLUUMI 2023-11-03 Turku Finland 176 -
Yi-34B 2023-11-02 01.AI China 34 1213
Skywork-13B 2023-10-30 Kunlun China 13 -
FinGPT-13B 2023-10-07 UCLA USA 13 -
Falcon-180B 2023-09-06 TII UAE 180 1034
Jais 2023-08-29 Cerebras Multi 13 -
Llama 2-70B 2023-07-18 Meta AI USA 70 1206
Llama 2-7B 2023-07-18 Meta AI USA - 1037
InternLM 2023-07-06 SAI Lab China 100 -
Goat-7B 2023-05-23 NUS Singapore 70 -
CodeT5+ 2023-05-20 Salesforce USA 160 -
CoEdiT-xxl 2023-05-17 Minnesota USA 110 -
PaLM 2 2023-05-10 Google USA 340 -
StarCoder 2023-05-09 Hugging Face Multi 155 -
Incoder-6.7B 2023-04-09 FAIR USA 67 -
BloombergGPT 2023-03-30 Bloomberg USA 505.588 -
Falcon-40B 2023-03-15 TII UAE 40 -
LLaMA-65B 2023-02-24 Meta AI USA 652 -
Hybrid H3-2.7B 2022-12-28 Stanford USA 27 -
GPT-3.5 Turbo 2022-11-30 OpenAI USA 200 1117
mT0-13B 2022-11-03 Hugging Face Multi 13 -
BLOOMZ-176B 2022-11-03 Hugging Face Multi 176 -
U-PaLM 2022-10-20 Google USA 540 -
LMSI-Palm 2022-10-20 Google USA 540 -
Flan-T5 11B 2022-10-20 Google USA 110 -
Flan-PaLM 2022-10-20 Google USA 540 -
BlenderBot 3 2022-08-10 McGill Canada 175 -
GLM-130B 2022-08-04 THU China 130 -
AlexaTM 20B 2022-08-02 Amazon USA 197.5 -
BLOOM-176B 2022-07-11 Hugging Face Multi 176 -
NLLB 2022-07-06 Meta AI USA 54.5 -
Minerva (540B) 2022-06-29 Google USA 540 -
UL2 2022-05-10 Google Multi 200 -
OPT-175B 2022-05-02 Meta AI USA 175 -
Sparse all-MLP 2022-04-14 Meta AI USA 94.1 -
PaLM (540B) 2022-04-04 Google Multi 540 -
Chinchilla 2022-03-29 DeepMind UK 70 -
DeepNet 2022-03-01 Microsoft USA 32 -
PolyCoder 2022-02-26 CMU USA 27 -
ST-MoE 2022-02-17 Google USA 269 -
LaMDA 2022-02-10 Google USA 137 -
GPT-NeoX-20B 2022-02-09 EleutherAI Multi 200 -
RETRO-7B 2022-02-07 DeepMind UK 75 -
AlphaCode 2022-02-02 DeepMind UK 411 -
InstructGPT 175B 2022-01-27 OpenAI USA 175 -
InstructGPT 6B 2022-01-27 OpenAI USA 60 -
InstructGPT 1.3B 2022-01-27 OpenAI USA 1.3 -
A.3 Discussion: Critical Reflections on Evaluation Practices
A.3.1 The Disconnect Between Evaluation Benchmarks and Real-World Performance
A critical challenge in contemporary LLM evaluation lies in the growing misalignment between standardized benchmarks and practical deployment requirements. While traditional evaluation methodologies provide valuable snapshots of model capabilities in controlled environments, they often fail to capture the nuanced interplay between model architecture, contextual adaptation, and real-world utility. This performance discrepancy reveals a fundamental limitation in current evaluation paradigms—their inability to adequately assess models in dynamic, interactive settings that better approximate production environments. The emergence of frameworks like Mint and WebArena represents a promising step toward addressing this gap by simulating realistic user interactions and environmental feedback loops, yet their adoption remains limited compared to traditional static benchmarks. This disconnect between laboratory evaluations and practical deployment outcomes has significant implications, as organizations increasingly rely on benchmark scores to make critical deployment decisions without fully understanding the limitations of these metrics in predicting real-world performance.
A.3.2 Fragmentation and Proliferation of Evaluation Benchmarks
The rapid proliferation of specialized evaluation benchmarks has created both opportunities and significant challenges for the research community. Analysis of numerous evaluation frameworks reveals substantial variation in model rankings across different benchmark categories, complicating cross-model comparison and creating what we term "evaluation overload." The situation is further exacerbated by the resource-intensive nature of comprehensive evaluation, which effectively excludes many academic and independent research groups from meaningful participation in rigorous model assessment. The knowledge base reveals an overwhelming diversity of benchmarks targeting specific capabilities, each with its own methodology and scoring system, making it difficult to synthesize a coherent understanding of model capabilities across the evaluation spectrum. This fragmentation hinders the development of a unified evaluation standard that could facilitate more meaningful progress in the field.
A.3.3 Language-Specific and Cultural Dimensions in LLM Evaluation
Evaluating language models in non-English contexts presents unique methodological challenges that extend beyond mere translation of English-centric benchmarks. The intricate nature of linguistic features in languages such as Chinese—including character-based semantics, tonal variations, and cultural context dependencies—requires specialized assessment frameworks that account for these distinctive characteristics. Current evaluation practices often overlook critical aspects such as idiomatic expression comprehension, classical language references, and culturally appropriate response generation. The knowledge base references several Chinese-specific evaluation frameworks like Zhujiu, yet these remain insufficient to address the full spectrum of linguistic and cultural nuances. This limitation extends beyond Chinese to numerous other languages, highlighting the urgent need for culturally adaptive metrics that assess not only linguistic accuracy but also sociocultural appropriateness within specific language contexts. The current evaluation ecosystem remains heavily biased toward English, with only a fraction of benchmarks addressing multilingual capabilities, thereby marginalizing the needs of the global majority of non-English language users.
A.3.4 Toward Integrated and Practical Evaluation Frameworks
Addressing the challenges outlined above requires the development of meta-evaluation frameworks that can synthesize results from multiple assessment dimensions while remaining accessible to resource-constrained researchers. Weighted aggregation approaches that prioritize benchmarks based on real-world task relevance rather than equal weighting offer a promising path forward, creating more meaningful composite scores that better predict practical model utility across diverse application scenarios. The knowledge base reveals several promising frameworks that could serve as building blocks for this integrated approach. An effective integrated evaluation framework should balance technical proficiency metrics (measured through standardized benchmarks), contextual adaptability (assessed via domain-specific tasks), and ethical robustness (evaluated through safety-oriented frameworks), creating a holistic assessment that better reflects real-world model performance.